A hand contains 2 red cards and 3 black cards. A card is selected from the hand and set aside. Then another card is selected and put with the first card. Let XX be the number of black cards drawn from the hand (i.e. set aside) during this process. Find Pr[X = 1]Pr[X=1].
Answer #1

For X = 1, either the first card should be black and second card red or the first card red and second card black

P(first card black and second card red) = 3/5*2/4 = 3/10

P(first card red and second card black) = 2/5*3/4 = 3/10

Thus, P(X = 1) = 3/10 + 3/10 = 0.6